OpenStudy (anonymous):

how do you write an equation in standard form for a line that passes throught the points -8, 4 and 4, -4?

OpenStudy (lgbasallote):

Two-point form: y - y1 = [(y2-y1)/(x2-x1)](x - x1) y -4 = [(-4-4)/(4 - -8)](x- -8) y - 4 = [(-8)/(12)](x+8) multiply both sides by 12 12y - 48 = -8x -64 simplifying further.... 3y - 12 = -2x - 16 so it;'s,... 2x + 3y + 4 =0